What is probability 2 girls are chosen if 10 students are girls and 15 are boys?
The probability of choosing 2 girls at random from group of 25 students of which
10 are girls and 15 are boys is:
P( 2 girls) = (10/25)∙(9/24) = 3/20 = 0.15 = 15%
(3/7)*(2/7)=(6/49)
You have a 6 out of 49 probability.
Number of possible ways to pull out one sheet = 8
Number of ways to pull out a number less than 5 = 4 (1, 2, 3, or 4)
Probability of success = 4/8 = 50%
What is the probabilit if I throw 4 dice probabiliti that I get all the same numbers on the dice?
(1/6) x (1/6) x (1/6) = 1/216 = 0.00463 = 0.463 percent
What is the definition of the Commutative Property of Addition?
Changing the order of addends does not change their sum. In symbolic form, a + b =b + a. Subtraction is not commutative .
Prob(No sixes in 6 rolls) = [Prob(Not a six in a roll)]6 = (5/6)6 = 0.3349
What information do you need to calculate a probability with a normal distribution?
Only the mean, because a normal distribution has a standard deviation equal to the square root of the mean.
The expected value is what kind of aspect of how probability distribution is characterized?
Expected value is the outcome of confidence of how probability distribution is characterized. If the expected value is greater than the confidence interval then the results are significant.
How many distinct ordered arrangements can be made with the letters of mississippi?
There are 34650 distinct orders.
There are 34650 distinct orders.
There are 34650 distinct orders.
There are 34650 distinct orders.
42 eyes
on a traditional deck of cards:
SPADES
jack-one
queen-two
king- two
DIAMONDS
jack- two
queen- two
king- one
CLUBS
jack- two
queen- two
king-two
HEARTS
jack- one
queen- two
king-two
What is the values of vitamin A to Z?
There is vitamin A (but no vitamin Z .. just the mineral Zinc).
How many ways can you rearrange the letters in the word turkey?
Since you didn't say they had to spell anything there are 720 possibilities.
What is the probabilty of throwing a 6 on a dice?
On a fair die, it is 1/6 on a single throw.
On a fair die, it is 1/6 on a single throw.
On a fair die, it is 1/6 on a single throw.
On a fair die, it is 1/6 on a single throw.
What is the probability of getting one prime number when rolling two dice?
In solving problems like this, the most straightforward approach is to make a table. Columns are used to show the six outcomes from one of the dice, rows the outcomes from the other die. In each cell of the table put the sum of the two dice. Now go throught the table row by row, counting the values that are prime numbers. There are 36 cells in the table, and I counted 15 primes, making a probability of 15/36 = 5/9.
But my counting has been wrong before! You had better check it.
Is 7k plus 0 equals 7k a multiplicative identity?
No because 0 is additive inverse the multiplicative inverse of a number is when you multiply that number by the m. inverse and still get that number.
How many cards are there in cardcaptors?
There is a total of 52 cards in Cardcaptors,like in a normal deck of cards. But, you don't get to see all 52 cards being captured. You only see 49 cards.
Benefits of probability to the business and administration?
one of the things i have learned from poker is that probability is not an indicator of what's to happen in the future, however, by playing the odds on a continuous basis you are assured to come out on top. just make sure to take the good odds, a 51/49 split is not a good chance in the world of probability. anything over 70% (i believe) is considered to be a good risk vs. reward assessment
40% chance
The term "8795896 heads" could refer to a numerical count of individuals or entities, such as people, animals, or any other countable items. Without additional context, it's difficult to determine the exact meaning or significance of this number. It could be used in various fields, such as demographics, livestock counting, or even metaphorically in discussions. If you have a specific context in mind, please provide more details for a more tailored response.
What is the definitions of a theoretical probability?
To me, the theoretical probability is what is termed the classical probability. This says the probability is the number of ways an event can occur divided by the number of possible events. Forexample, flip a coin. The theoretical probability for heads is 1/2. However, flip a coin 10 times and you will probably not get 5/10 (or 1/2). Doing the actual experiment to determine the probability is called relative frequency approximation.
A die is rolled and a coin is tossed how many have heads as the coin result?
There are 6 rolls for the die times 2 possibilities for the coin, so there are 12 total possibilities. Because there is a 1/2 chance of flipping a head, the answer is half of 12, or 6.
What is your grade if you miss 12 out of 43?
If you missed 12, then you got 31 correct [43-12=31]. If all questions are weighted equally, then 31/43 = 0.721 or 72.1 %, so your grade is 72.